The following is the cost of an article at a capacity level of 5000 units as...

  1. The following is the cost of an article at a capacity level of 5000 units as given under A. For a variation of 25% in capacity above or below this level, the individual expenses vary as indicated under B.

                                                                     A                            B                                                    

                                                      $                             %

Material Cost                              25,000                    100

Labor Cost                                   15,000                    100

Power                                             1,250                      80

Repairs and maintenance           2,000                      75

Stores                                             1,000                    100

Inspection                                        500                      20

Depreciation                               10,000                        -

Administrative OH                        5,000                     25

Selling                                              3,000                      50

Total Cost                                     62,750

Cost per unit                                                 12.55

Required:

Find the unit cost of production under each individual expense at the level of 4000 and 6000 units by preparing a flexible budget.
Answer #1

Fixed costs are those costs which would remain the same in total irrespective of level of production or sales. It is a commitAt 6,000 units Cost per Units Fixed costs Total costs Variable Cost per unit unit (e)/(b) (b) (d) (e) = (c) + (d) (a) $0 TotaWorkings: C=AB D=A-C E Dollar Variable % Variable $ Fixed $ Units 100% 100% 80% $25,000 $15,000 $1,000 $1,500 $1,000 Material